While playing at Azadi Stadium, Ali threw a soccer ball that reached a maximum height of 30 meters. If he did not defend, he would have landed at a distance of 70 meters. Yes, Ali throws the ball when the ball is 2 meters above the ground and there is no air resistance, at what horizontal distance does the defense catch the ball?

1
Expert's answer
2021-11-05T18:19:11-0400

distance:

"x=v_{0x}t"

height:

"h=v_{0y}t-gt^2\/2"


maximum height:

"h'(t)=v_{0y}-gt=0"


"h_{max}=h_0+v_{0y}^2\/g-v_{0y}^2\/2g=30" m


"v_{0y}=\sqrt{2gh_{max}}=\sqrt{2\cdot 9.8\cdot30}=24.25" m/s


time of maximum height:

"t_{m}=v_{0y}\/g=24.25\/9.8=2.47" s


distance:

"x=v_{0x}t=2v_{0x}t_m=70" m


"v_{0x}=x/(2t_m)=70/(2\cdot 2.47)=14.17" m/s


if Ali throws the ball when the ball is 2 meters above the ground:


"h=2+v_{0y}t-gt^2\/2"


for h=0:


"4.9t^2-24.25t-2=0"


"t=\frac{24.25+ \sqrt{24.25^2+8\cdot4.9}}{9.8}=5.03" s


So:


"x=14.17\cdot5.03=71.28" m